OpenStudy (anonymous):

Simplify each expression. Rationalize all denominators. Assume that all variables are positive. 1) 1 over square root of 3 + 5 (it's 1 over the whole thing, so it's 1 over square root of 3 and the + 5 isn't under the radical but it's still in the denominator) 2) square root of 63 + 2 square root of 28 - 5 square root of 7 Please explain and show steps!

OpenStudy (anonymous):

\[1/\sqrt{3} + 5\]

OpenStudy (anonymous):

@imranmeah91 @ishaan94 @pokemon23 someone help

OpenStudy (pokemon23):

Ishaan got this.. pro math wizard

OpenStudy (anonymous):

\[\frac{\sqrt{3} - 5}{\sqrt{3} - 5}\times \frac{1}{\sqrt{3} + 5}\] \[\frac{\sqrt{3} - 5}{3 - 5} = \frac{\sqrt{3} - 5}{-2} = \frac{5 - \sqrt{3}}{2}\]

OpenStudy (pokemon23):

like a boss

OpenStudy (anonymous):

lol, like a sir

OpenStudy (anonymous):

@trolololcat Can you do the second part on your own now? Or, need I show you second one too?

OpenStudy (anonymous):

This is really new to me show me the second one too please

OpenStudy (anonymous):

2. square root of 63 + 2 square root of 28 - 5 square root of 7 Solution: \(\sqrt{63} + 2 \sqrt{28} - 5 \sqrt{7} = \sqrt{7 \times 9} + 2 \sqrt{4 \times 7} - 5\sqrt{7} \) \[\implies\sqrt{7 \times 3^2} + 2 \sqrt{2^2 \times 7} - 5\sqrt{7} = 3\sqrt{7} + 4\sqrt{7} - 5\sqrt{7} \] \[\implies \sqrt{7} \left(3 + 4 - 5 \right) = 2\sqrt{7} \]

OpenStudy (anonymous):

@ishaan94 Do you think you can help me on two more questions? >.< 1)\[\sqrt[3]{2}+1/\sqrt[3]{4}\] 2) 2/1+ square root of 2

OpenStudy (anonymous):

I hope you meant to simply the expression, right?

OpenStudy (anonymous):

\[1. \:\:\:\:\:\: \ \frac{\sqrt[3]{2} + 1}{\sqrt[3]{2^2}} \implies \large \frac{2^{\frac{1}{3}}}{2^{\frac{2}{3}}} + \frac{1}{\sqrt[3]{2^2}} \implies 2^{\frac{-1}{3}} + \frac{1}{\sqrt[3]{2^2}}\]

OpenStudy (anonymous):

yes to simplify and rationalize denominators

OpenStudy (anonymous):

I'm confused about 1), what did you do first and what exactly is the answer o_o

OpenStudy (anonymous):

argh, I'm sorry... It took time I was on chat :/ Lets do it again, ignore my previous solution... \[\frac{\sqrt[3]{2} + 1}{\sqrt[3]{2^2}} \times \frac{\sqrt[3]{2}}{\sqrt[3]{2}} = \frac{\sqrt[3]{4} + \sqrt[3]{2}}{2}\]

OpenStudy (anonymous):

What I did is got rid of the cube-root of the denominator, that's it, that's all I have done by multiplying and dividing \(\sqrt[3]{2}\).

OpenStudy (anonymous):

For the second part. \[\frac{1}{1 + \sqrt{2}} \times \frac{1-\sqrt{2}}{1 - \sqrt{2}} = \frac{1 - \sqrt{2}}{1 - 2} = \frac{\sqrt{2} -1}{1} = \sqrt{2} - 1\]

OpenStudy (anonymous):

so its \[\sqrt[3]{6}/2 ?\]

OpenStudy (anonymous):

No, for the first question, I don't think you can simplify it further than my final expression :/

OpenStudy (anonymous):

@ishaan94 why can't you combine them?

OpenStudy (anonymous):

@Chlorophyll can you double check 1) for us of the second set lol

OpenStudy (anonymous):

Um their powers are different, when powers are different you can multiply them to each-other, divide them but can not add or subtract.

OpenStudy (anonymous):

@Ishaan94 dont both of them have the power of 3?

OpenStudy (anonymous):

argh what I meant was if both the bases and the powers are same, you need both the base and power to be equal to add or subtract.

OpenStudy (anonymous):

@Ishaan94 so you think its safe to put that as my final answer

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

@amistre64 you think so too?

OpenStudy (amistre64):

im reading the original post and seeing something different get answered

OpenStudy (amistre64):

what is the original problem?

OpenStudy (amistre64):

1) 1 over square root of 3 + 5 (it's 1 over the whole thing, so it's 1 over square root of 3 and the + 5 isn't under the radical but it's still in the denominator) 1/(sqrt(3) + 5)

OpenStudy (amistre64):

then yeah, i read the original wrong

OpenStudy (amistre64):

a^2-b^2 = (a+b)(a-b) so if we have somehting of the form (a-b) we multiply it by an opposite operation to rationalize it

OpenStudy (anonymous):

amistre, this one ↓ ↓\[\frac{\sqrt[3]{2} + 1}{\sqrt[3]{2^2}} \times \frac{\sqrt[3]{2}}{\sqrt[3]{2}} = \frac{\sqrt[3]{4} + \sqrt[3]{2}}{2}\] Can you simplify it further? or, Add the terms in Numerator? I don't think so.

OpenStudy (amistre64):

you can factor out a cbrt(2) but then you created it to begin so thats not realy a simplification is it otherwise it loks fine to me

OpenStudy (anonymous):

It's probably correct by applying conjugation: 1/ (sqrt3 + 5) = 1/ (sqrt3 + 5) * (sqrt3 - 5)/ (sqrt3 -5) = ( sqrt3 -5) / ( 3 - 25) = -( sqrt3 -5) / 22

OpenStudy (anonymous):

My PC's too slow at peak time!

OpenStudy (anonymous):

= ( 5- sqrt3 ) / 22 Look much more beautiful :)